OpenStudy (anonymous):

\[\sqrt[3]{x}^{2} \times \sqrt[4]{x}\] \[\sqrt[3]{x}^{2} \times \sqrt[4]{x}\] @Mathematics

OpenStudy (anonymous):

x^(2/3) x x^(1/4) x^(2/3+1/4) x^ (8/12 + 3/12) =x^(11/12) \[\sqrt[12]{x}^{11}\]